-4/9+ 3/10
. Perform the indicated operation and, if possible, simplify. Please show all of your work

Answer by vasumathi(46)

-4/9+ 3/10 To add any two fractions we need to make the denominators common first
multiply both numerator as well as the denominator of the first one by 10 and ...
multiply both numerator as well as the denominator of the second one by 9.We have ...
-4*10/9*10 +3*9/10*9=-40/90+27/90= (-40+27 )/90=-13/90 This is the final answer